# LLMs as Span Annotators: A Comparative Study of LLMs and Humans

### v3

@@ 7 Conclusion
We showed that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Automating span annotation with LLMs seems to be a promising alternative to fine-grained human evaluation sourced from crowdsourcing platforms, opening the way towards scalable and actionable automatic NLG evaluation methods.
@@
Although we aimed to select a representative sample of models, prompts and tasks, our choice is constrained by our limited time frame and budget. Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their ambiguity, the annotation guidelines, or the qualification level of human annotators. The estimates are also not readily available for existing datasets and require additional data collection.
@@ Acknowledgments
This work was funded by the European Union (ERC, NG-NLG, 101039303) and the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
We thank David M. Howcroft for his early input and contributions to the research methodologies adopted in this study.
@@ A.1 Open Models
We run the local models using the ollama framework in 4-bit quantization. Specifically, we use llama3.3:70b and deepseek-r1:70b for span annotations. We also use gemma2:2b and phi3.5:3.8b for generating texts in D2T-Eval.
@@ A.3 Web Interface
We implement our span annotation process using factgenie kasner2024factgenie: a tool that supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
@@ B.1 Generating Outputs
For generating the outputs for the structured inputs we collected, we use two larger models – Llama 3.3 70B grattafiori2024llama and GPT-4o hurst2024gpt – and two smaller models – Gemma 2 2B team2024gemma and Phi-3.5 3.8B abdin2024phi. See more details on the models in Appendix˜A and prompts in Appendix˜D.
@@ B.2 Collecting annotations
For the annotation guidelines, we went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works kasner2024traditional; thomson2020gold. We settled on the following annotation categories (see Table˜5 for details): semantic accuracy errors due to information Contradictory to the input, Not Checkable, or Misleading; any Incoherent and Repetitive content, and any Other errors.
@@
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
@@ Appendix C Annotation Categories
Table 7: Annotation categories for the Propaganda task. The categories are adopted from dasanmartino2019finegrained.

### v4

@@ 7 Conclusion
We showed that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to verified crowdworkers who passed a qualification task. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Automating span annotation with LLMs seems to be a promising alternative to fine-grained human evaluation sourced from crowdsourcing platforms, opening the way towards scalable and actionable automatic NLG evaluation methods.
@@
Although we aimed to select a representative sample of models, prompts and tasks, our choice is constrained by our limited time frame and budget. Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their ambiguity, the annotation guidelines, or the qualification level of human annotators. The estimates are also not readily available for existing datasets and require additional data collection. Due to our insufficient expertise in the target languages, we also do not provide language-specific manual error analysis of results.
@@ Acknowledgments
This work was funded by the European Union (ERC, NG-NLG, 101039303). It was additionally supported by the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO and the Charles University Research Centre program No. 24/SSH/009. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2023062). We thank David M. Howcroft for his early input and contributions to the research methodologies adopted in this study.
@@ A.1 Open Models
We run the local models using the ollama framework in 4-bit quantization. Specifically, we use llama3.3:70b and deepseek-r1:70b (which is based on Llama 3.3 70B) for span annotations. We also use gemma2:2b and phi3.5:3.8b for generating texts in D2T-Eval.
@@ A.3 Web Interface
We implement our span annotation process using factgenie Kasner et al. (2024): a tool that supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
@@ B.1 Generating Outputs
For generating the outputs for the structured inputs we collected, we use two larger models – Llama 3.3 70B Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024) – and two smaller models – Gemma 2 2B Team et al. (2024) and Phi-3.5 3.8B Abdin et al. (2024). See more details on the models in Appendix˜A and prompts in Appendix˜D.
@@ B.2 Collecting annotations
For the annotation guidelines, we went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner and Dušek (2024); Thomson and Reiter (2020). We settled on the following annotation categories (see Table˜5 for details): semantic accuracy errors due to information Contradictory to the input, Not Checkable, or Misleading; any Incoherent and Repetitive content, and any Other errors.
@@
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. Our process of collecting annotations proceeded in two stages, following the setup of Zhang et al. (2023): (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
@@ Appendix C Annotation Categories
Table 7: Annotation categories for the Propaganda task. The categories are adopted from Da San Martino et al. (2019).